4x+3y=-12 explain please

Answer by lynnlo(4176)

PROBLEM:
4X+3Y=-12
ADD "-3Y" TO EACH SIDE OF EQUATION
4X+3Y+-3Y=-12+-3Y
COMBINE LIKE TERMS 3Y+-3Y=0
4X+0=-12+-3Y
4X=-12+-3Y
DIVIDE EACH SIDE BY "4"
X=-3+-0.75Y
SIMPLIFYING
X=-3+-0.75Y
ANSWER
X=-3Y/4-3